### v2

@@ Abstract
Expert parallelism has been introduced as a strategy to distribute the computational workload of sparsely-gated mixture-of-experts (MoE) models across multiple computing devices, facilitating the execution of these increasingly large-scale models. However, the All-to-All communication intrinsic to expert parallelism constitutes a significant overhead, diminishing the MoE models’ efficiency. Current optimization approaches offer some relief, yet they are constrained by the sequential interdependence of communication and computation operations. To address this limitation, we present a novel shortcut-connected MoE (ScMoE) architecture with an overlapping parallel strategy, which effectively decouples communication from its conventional sequence, allowing for a substantial overlap of 70% to 100% with computation. When compared with the prevalent top-2 MoE architecture, ScMoE demonstrates training speed improvements of 30% and 11%, and inference improvements of 40% and 15%, in our distributed environments with PCIe and NVLink hardware, respectively, where communication constitutes 60% and 15% of the total MoE time consumption. Building on the ScMoE architecture, we further implement an expert offloading strategy to facilitate memory-limited inference, optimizing latency through the overlap of expert migration. Additionally, extensive experiments and theoretical analyses indicate that ScMoE not only achieves comparable but in some instances surpasses the model quality of existing approaches.
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
@@ 1 Introduction
In recent years, Transformer-based large language models (LLMs) have significantly propelled the fields of Natural Language Processing Vaswani et al. (2017); Brown et al. (2020); Wei et al. (2022b); Ouyang et al. (2022); Wei et al. (2022a); Chowdhery et al. (2023); Achiam et al. (2023), Computer Vision Dosovitskiy et al. (2021); Liu et al. (2021), and Multimodality Lu et al. (2019); Zhou et al. (2022); Zhang et al. (2021); Zhu et al. (2023). The sparsely-gated mixture-of-experts (MoE) approach has been integral in increasing parameter counts and enhancing model performance across various modalities Shazeer et al. (2017); Riquelme et al. (2021); Mustafa et al. (2022); Jiang et al. (2024). Expert parallelism Lepikhin et al. (2021); Fedus et al. (2022) has emerged as a viable strategy to distribute MoE computations efficiently over multiple devices, synergizing with conventional parallelism techniques Hwang et al. (2022); Singh et al. (2023) such as data parallelism Rajbhandari et al. (2020; 2021) and model parallelism Narayanan et al. (2021); Smith et al. (2022).
@@
Nevertheless, expert parallelism incurs substantial All-to-All communication overhead Lepikhin et al. (2021); Fedus et al. (2022), which can contribute to approximately 50% of the total time in intra-node multi-GPUs or multi-nodes distributed environments (see Figure 1), thus forming a critical bottleneck in scaling MoE models Nie et al. (2022); Hwang et al. (2022); Mayer & Jacobsen (2020); Smith et al. (2022). Despite existing optimizations such as hierarchical All-to-All He et al. (2022); Nie et al. (2022) and pipelining Hwang et al. (2022); Zhang et al. (2023) strategies that mitigate communication delays and partially overlap communication with computation, the communication challenge persists due to the inherent sequential dependencies between these operations Wang et al. (2023). To address this constraint, our intuitive idea is to reconstruct the inputs of MoE layer by incorporating not only the current-layer but also the preceding-layer representations through a shortcut connection, thereby refining the communication dependencies and expanding the potential for optimization.
@@
Figure 1: The overhead of MLP and top-2/top-1 MoE in a transformer block of SwinV2-MoE-S Hwang et al. (2022) model, allocating one expert per GPU with expert parallelism. The All-to-All communication takes up 60% of total time on a single node with 8$\times$×A30 GPUs, but drops to 15% on 8$\times$×A800 due to the latter’s 6$\times$ higher bandwidth provided by GPU-to-GPU NVLink Foley & Danskin (2017). Despite benefiting from NVLink, communication still approaches 50% due to the lower-bandwidth inter-node Ethernet Li et al. (2020) when scaling across multiple nodes.
@@
In this paper, we introduce the shortcut-connected MoE (ScMoE) architecture, which completely decouples communication processes from the sequence of conventional MoE models. Our architecture is built upon the standard top-2 MoE (see Figure 2 (a)) which typically substitutes the Multi-Layer Perceptron (MLP) module with a top-2 gating MoE module in every second Transformer block (refer to the Transformer block with MoE module as “current layer”, and the preceding one without MoE module as “preceding layer”). Diverging from the top-2 approach, our ScMoE utilizes a top-1 MoE module to process preceding-layer representations via a shortcut connection, while employing a shared expert (an MLP module) to process current-layer representations. These two processes are independently managed in parallel, with their results integrated into the final output of the current layer. Furthermore, we select the intermediate representations between the Attention and MLP modules in the preceding layer as the input for the top-1 MoE. Our experiments on various preceding-layer representations (output, intermediate, input) indicate that this choice achieves optimal accuracy and a relatively longer overlap duration.
@@
To efficiently overlap the decoupled communication and computation within our ScMoE architecture, we implement an adaptive overlapping parallel strategy that dynamically schedules operators based on actual performance metrics. Relative to existing optimization strategies, our approach not only doubles the overlap duration compared to the pipelining Huang et al. (2019); Hwang et al. (2022) but also realizes complete overlapping of communication in scenarios where communication time does not surpass the computation duration. Furthermore, our method essentially advances the MoE architecture in algorithm aspect, which is device-agnostic to improve the efficiency of MoE model, thus ensuring a broad applicability across various hardware configurations and maintaining compatibility with current optimizations.
Additionally, existing studies Hwang et al. (2023b); Yi et al. (2023) offload expert parameters to CPU memory in memory-limited inference scenarios where GPU cannot store the full MoE model. These studies utilize information in preceding layers to predict expert selection for the current MoE layer, enabling early expert migration from CPU to GPU and overlapping it with model computation. In contrast to existing speculative expert migration methods, we implement an expert offloading strategy with overlapping determinate migration, built upon our ScMoE that inherently advances expert selection to the preceding layer.
@@
The extensive experimental results reveal that compared with the standard top-2 MoE, our ScMoE architecture achieves training speed improvements of 30% and 11% in 8$\times$×A30-PCIe and 8$\times$×A800-NVLink scenarios characterized by high and low communication overheads, respectively, and deliver inference speed improvements of 40% and 15%. Besides, our ScMoE has been demonstrated through experiments and theoretical analysis to attain or exceed the model quality of existing methods in both vision and language downstream tasks. In memory-limited inference, our expert offloading strategy reduces peak GPU memory usage by up to 60% and decreases expert migration costs by up to 75% through overlapping with computation. In addition, we conduct an in-depth analysis of our ScMoE, investigating the proposed shortcut connection and exploring the potential reasons for its algorithmic effectiveness.
@@
  • We propose the shortcut-connected MoE (ScMoE) architecture that breaks conventional interdependency between communication and computation in distributed MoE models, bypassing the restrictions imposed on current communication optimization techniques.
  • We develop an adaptive overlapping strategy for advancing expert parallelism with our shortcut-connected MoE, which significantly improves the efficiency of MoE models and ensures broad compatibility.
  • We implement an expert offloading strategy with overlapping determinate expert migration upon our ScMoE, enhancing efficiency in memory-limited inference.
@@ 2.1 Sparsely-Gated Mixture of Experts
The Sparsely-Gated Mixture-of-Experts Shazeer et al. (2017) (MoE) layer is composed of multiple Multi-Layer Perceptron (MLP) sub-networks, termed “experts”, and employs a trainable gating network to selectively activate a subset of these experts during each iteration. Given $N$ expert networks $\{E_{i}\}_{1}^{N}$, gating network $G$ and input representation $x$, the output of MoE module can be written:
$MoE(x)=\sum_{i=1}^{N}G(x)_{i}E_{i}(x)$ (1)
@@
$G(x)=Softmax(\overline{TopK}(H(x),k))$ (2)
@@
Leveraging sparse output of $G(x)$, this approach significantly increases the number of model parameters without causing a proportional increase in computational demand. The value of $k$ can be set to 1 or 2 or even higher values. Opting for a larger $k$ moves the model closer to the dense architecture, which generally results in higher prediction accuracy Riquelme et al. (2021), but also leads to larger execution overhead.
@@
Figure 2 (a) illustrates the prevailing top-2 MoE architecture. Each Transformer block with MoE module, denoted by a light blue block and referred to as “Block-MoE,” replaces the MLP with a set of experts (“E1, E2, E3, E4”) and a gating network (“Gate”). Following prior work Lepikhin et al. (2021); Du et al. (2022); Shen et al. (2023); Hwang et al. (2022), the “Block-MoE” is interspersed with the conventional Transformer block, depicted as a gray block and referred to as “Block-MLP.”
@@
$MoE(x)=SE(x)+\sum_{i=1}^{N}G(x)_{i}E_{i}(x)$ (6)
This method, initially proposed by DeepSpeed-MoE Rajbhandari et al. (2022), activates the same number of experts as standard MoE for computation while reducing dynamic expert selection and communication volume. Extensive empirical results from DeepSpeed-MoE and subsequent studies Dai et al. (2024); Team (2024) demonstrate that the shared expert achieves on-par or even better model quality,leading to its increasing application Xue et al. (2024); Wu et al. (2023); Chen et al. (2024); Gou et al. (2023); Gao et al. (2024).
@@ 2.2 Expert Parallelism
Figure 4: Illustrations of various ScMoE architectures with shortcut connections to different positions of the preceding layer: (a) “Pos-1” output, (b) “Pos-2” intermediate, and (c) “Pos-3” input. The red line indicates the transmission of the preceding-layer representations to the MoE via a shortcut connection. To streamline the explanation, details regarding pre-layer normalization (Pre-Norm) and dropout procedures have been excluded. The ScMoE (Pos-2) configuration (b) is selected as the default ScMoE architecture for further implementations and experiments due to its superior performance, as shown in Table 1.
To facilitate efficient distributed training and inference of MoE models, expert parallelism is proposed to allocate unique experts to each distributed computing device such as GPU and TPU, and map tokens to their corresponding experts through All-to-All communication across the participating devices Lepikhin et al. (2021); He et al. (2021); Nie et al. (2022). As illustrated in Figure 3, the workflow of MoE employing expert parallelism is segmented into the following sequential operations: gate routing, input encode, All-to-All dispatch, expert computation, All-to-All combine, and output decode. To enhance the efficiency, input encode is employed to aggregate the token data layout to a contiguous format before All-to-All dispatch, and output decode is the inverse process after All-to-All combine. Furthermore, the integration of expert parallelism with other parallelisms Hwang et al. (2022); Singh et al. (2023); Fedus et al. (2022); Zheng et al. (2022) has been explored to support the scaling of larger MoE models on extensive distributed systems.
However, the All-to-All communication used for token transfer has been a primary factor limiting the efficiency of distributed MoE models, as illustrated in Fig 1. To address this issue, we enhance the MoE architecture and expert parallelism with shortcut connections, significantly overlapping communication costs with computation.
2.3 Memory-Limited Inference of MoE Models
While MoE effectively enhances LLMs in terms of model quality, it faces significant deployment challenges during on-device inference due to high memory demand. A common approach is to offload expert parameters to CPU memory Shen et al. (2022) in scenarios where GPU memory is insufficient to store the entire MoE model. Since the migration of activated expert parameters from CPU to GPU, which occurs after expert selection, blocks expert computation until the transfer is complete, existing studies Hwang et al. (2023b); Yi et al. (2023); Du et al. (2024) have explored prefetching the experts. For instance, Pre-gated MoE Hwang et al. (2023b) uses information from preceding layers to predict expert selection, allowing for preloading of expert parameters into GPU memory, as shown in Figure 7. This method enables overlapping the expert migration duration with the computation of preceding modules.
However, speculative expert migrations can suffer from estimation inaccuracies, as they deviate from the original logic of pre-trained models, potentially reducing inference accuracy. In contrast, we enhance the MoE architecture with a shortcut connection, utilizing the MoE module to process preceding-layer representations. This allows us to implement an expert offloading strategy with overlapping determinate migration, maintaining the pre-trained logic.
@@ 3 Shortcut-connected MoE Designs
In the prevailing Transformer-based model, the MoE module substitutes MLP to sequentially manipulate intermediate representations Lepikhin et al. (2021); Du et al. (2022); Shen et al. (2023), impeding the efficacy of existing optimization strategies He et al. (2022); Nie et al. (2022); Hwang et al. (2022); Zhang et al. (2023) due to the limited interaction within the MoE module. To address the aforementioned limitations, we propose the shortcut-connected MoE (ScMoE) architecture (Section 3.1), which enables optimization opportunities for computation-communication overlap for expert parallelism (Section 3.2) and determinate expert migration for memory-limited inference (Section 3.3).
Table 1: Test top-1 accuracy and overlap duration of pre-trained SwinV2-MoE-S models with various ScMoE configurations (illustrated in Figure 4). $\mathcal{T}_{Atten}$, $\mathcal{T}_{SE}$, and $\mathcal{T}_{MLP}$ represent the durations of Attention, shared expert, and MLP, respectively.
ScMoE ImageNet-1K (Acc@1$\uparrow$) Overlap Duration
Pos-1 79.14% $\mathcal{T}_{Atten}+\mathcal{T}_{SE}$
Pos-2 79.38% $\mathcal{T}_{Atten}+\mathcal{T}_{SE}+\mathcal{T}_{MLP}$
Pos-3 79.20% $2\mathcal{T}_{Atten}+\mathcal{T}_{SE}+\mathcal{T}_{MLP}$
@@ 3.1 Architectural Design
In this section, we introduce the shortcut-connected MoE (ScMoE) architecture. Unlike prevailing MoE architectures, illustrated in Figure 2, which focus on processing intermediate representations within the current layer (the Transformer block containing the MoE), our ScMoE processes representations from both the current and preceding layers. Specifically, ScMoE employs a top-1 MoE module to handle representations from the preceding layer via a shortcut connection, while a shared expert processes the current-layer representations. These two operations are conducted independently, with their outcomes integrated into the final output of the current layer, facilitating communication and computation overlap between these two processes.
While the shared expert processes the same intermediate representations in the current layer as the prevailing MoE approaches, we explore three distinct preceding-layer representations for ScMoE’s top-1 MoE process, as illustrated in Figure 4. The configurations “Pos-1” (a), “Pos-2” (b), and “Pos-3” (c) represent shortcuts connecting different positions of the preceding layer: output, intermediate, and input, respectively.
Based on our experimental results presented in Table 1, the configuration “Pos-2,” which selects intermediate representations between the Attention and MLP modules in the preceding layer as input for the top-1 MoE module, achieves optimal accuracy (79.38%) and the second longest overlap duration ($\mathcal{T}_{Atten}+\mathcal{T}_{SE}+\mathcal{T}_{MLP}$). Consequently, we have chosen ScMoE (Pos-2) as the default ScMoE architecture for subsequent implementations and experiments. This ScMoE architecture can be formulated as follows:
@@
$+\sum_{i=1}^{N}G(\mathcal{H}^{MH}_{l})_{i}E_{i}(\mathcal{H}^{MH}_{l})$
$\mathcal{H}^{MH}_{l+1}=\mathcal{H}^{MLP}_{l}+\text{MultiHead}^{(l+1)}_{MoE}(\mathcal{H}^{MLP}_{l})$ (8)
@@
$\mathcal{H}^{MLP}_{l}=\mathcal{H}^{MH}_{l}+\text{MLP}^{(l)}(\mathcal{H}^{MH}_{l})$ (9)
$\mathcal{H}^{MH}_{l}=\mathcal{H}_{l-1}+\text{MultiHead}^{(l)}_{MLP}(\mathcal{H}_{l-1})$ (10)
where $\mathcal{H}_{l+1}^{\text{ScMoE}}$ refers to the output from the MoE sub-layer, $\mathcal{H}_{l+1}^{MH}$ signifies the output from the Multi-Head Attention (MultiHead) sub-layer $\text{MultiHead}^{(l+1)}_{MoE}(\cdot)$ in the $(l+1)$-th Transformer block (“Block-MoE”). $SE^{(l+1)}(\cdot)$ denotes the shared expert while $E_{1},...,E_{n}$ represent the $N$ gate-selected experts. The gating network $G(\cdot)$ is referred to as Equation 2. $\mathcal{H}_{l}^{MLP}$ and $\mathcal{H}_{l}^{MH}$ are the outputs of the MLP sub-layer $\text{MLP}^{(l)}(\cdot)$ and the MultiHead sub-layer $\text{MultiHead}^{(l)}_{MLP}(\cdot)$, respectively, in the $l$-th Transformer block (“Block-MLP”). Note that we omit the pre-layer normalization and dropout for simplicity.
In addition, we provide an in-depth theoretical analysis of our proposed ScMoE architecture in Appendix A.1, elucidating the propagation of gradients to guarantee consistent training and preserve model quality.

### v3

@@ Abstract
Expert parallelism has emerged as a key strategy for distributing the computational workload of sparsely-gated mixture-of-experts (MoE) models across multiple devices, enabling the processing of increasingly large-scale models. However, the All-to-All communication inherent to expert parallelism poses a significant bottleneck, limiting the efficiency of MoE models. Although existing optimization methods partially mitigate this issue, they remain constrained by the sequential dependency between communication and computation operations. To address this challenge, we propose ScMoE, a novel shortcut-connected MoE architecture integrated with an overlapping parallelization strategy. ScMoE decouples communication from its conventional sequential ordering, enabling up to 100% overlap with computation. Compared to the prevalent top-2 MoE baseline, ScMoE achieves speedups of 1.49×1.49$\times$1.49 × in training and 1.82×1.82$\times$1.82 × in inference. Moreover, our experiments and analyses indicate that ScMoE not only achieves comparable but in some instances surpasses the model quality of existing approaches.
Mixture of experts, Expert parallelism, Shortcut connection
@@ 1 Introduction
In recent years, Transformer-based large language models (LLMs) have significantly propelled the fields of Natural Language Processing (Vaswani et al., 2017; Brown et al., 2020; Wei et al., 2022b; Ouyang et al., 2022; Wei et al., 2022a; Chowdhery et al., 2023; Achiam et al., 2023), Computer Vision (Dosovitskiy et al., 2021; Liu et al., 2021), and Multimodality (Lu et al., 2019; Zhou et al., 2022; Zhang et al., 2021; Zhu et al., 2023). The sparsely-gated mixture-of-experts (MoE) approach has been integral in increasing parameter counts and enhancing model performance across various modalities (Shazeer et al., 2017; Riquelme et al., 2021; Mustafa et al., 2022; Jiang et al., 2024). Expert parallelism (Lepikhin et al., 2021; Fedus et al., 2022) has emerged as a viable strategy to efficiently distribute MoE computations over multiple devices, synergizing with conventional parallelism techniques (Hwang et al., 2023; Singh et al., 2023) such as data parallelism (Rajbhandari et al., 2020, 2021) and model parallelism (Narayanan et al., 2021; Smith et al., 2022).
@@
Nevertheless, expert parallelism incurs substantial All-to-All communication overhead (Lepikhin et al., 2021; Fedus et al., 2022), which can contribute to approximately 50% of the total time in intra-node multi-GPUs or multi-nodes distributed environments (see Figure 1), thus forming a critical bottleneck in scaling MoE models (Nie et al., 2022; Hwang et al., 2023; Mayer & Jacobsen, 2020; Smith et al., 2022). Despite existing optimizations such as hierarchical All-to-All (He et al., 2022; Nie et al., 2022) and pipelining (Hwang et al., 2023; Zhang et al., 2023) strategies that mitigate communication delays and partially overlap communication with computation, the communication challenge persists due to the inherent sequential dependencies between these operations (Wang et al., 2023). To address this constraint, our intuitive idea is to reconstruct the inputs of MoE layer by incorporating not only the current-layer but also the preceding-layer representations through a shortcut connection, thereby refining the communication-computation dependencies and expanding the potential for their overlapping optimization.
@@
Figure 1: The overhead of MLP and top-2/top-1 MoE in a transformer block of SwinV2-MoE-S (Hwang et al., 2023) model, allocating one expert per GPU with expert parallelism. The All-to-All communication takes up 60% of total time on a single node with 8$\times$×A30 GPUs, but drops to 15% on 8$\times$×A800 due to the latter’s 6$\times$ higher bandwidth provided by GPU-to-GPU NVLink (Foley & Danskin, 2017). Despite benefiting from NVLink, communication still approaches 50% due to the lower-bandwidth inter-node Ethernet (Li et al., 2020) when scaling across multiple nodes.
@@
In this paper, we propose the shortcut-connected MoE (ScMoE) architecture, which completely decouples communication processes from the sequence of conventional MoE models. ScMoE architecture is initially built on the standard top-2 MoE (see Figure 2 (a)), which typically substitutes the multi-layer perceptron (MLP) module with a top-2 gating MoE module in every second Transformer block (refer to the Transformer block with the MoE module as “current layer”, and the preceding one without the MoE module as “preceding layer”). Diverging from the top-2 approach, ScMoE utilizes a top-1 MoE module to process preceding-layer representations via a shortcut connection, while employing a shared expert (an MLP module) to process current-layer representations. These two processes are independently managed in parallel, with their results integrated into the final output of the current layer. Furthermore, the ScMoE architecture can be extended to accommodate MoE models that employ various MoE placement frequencies, such as integrating an MoE module into every Transformer block.
@@
To efficiently overlap the decoupled communication and computation within the ScMoE architecture, we implement an adaptive overlapping parallel strategy that dynamically schedules operators based on actual performance metrics. Compared to existing optimization strategies such as pipelining (Huang et al., 2019; Hwang et al., 2023), our approach not only doubles the overlap duration, but also realizes complete overlapping of communication in scenarios where communication time does not exceed the computation duration. Furthermore, our method essentially advances the MoE architecture in algorithm aspect, which is device-agnostic to improve the efficiency of MoE model, thus ensuring a broad applicability across various hardware configurations and maintaining compatibility with current optimizations.
@@
The experimental results reveal that, compared to the standard top-2 MoE, our proposed ScMoE architecture optimally accelerates training by 1.49×1.49$\times$1.49 × and 1.14×1.14$\times$1.14 × in 8$\times$×A30-PCIe and 8$\times$×A800-NVLink scenarios characterized by high and low communication overheads, respectively, and accelerates inference by 1.82×1.82$\times$1.82 × and 1.21×1.21$\times$1.21 ×. Moreover, we perform experiments on different configurations of the ScMoE architecture, including shortcut-connected position and coefficient gating network. Considering the optimal accuracy and the relatively longer overlap duration, we favor selecting the intermediate representations between the Attention and MLP modules in the preceding layer as the input for the gate-routed experts. In addition, ScMoE has been demonstrated through experiments and theoretical analysis to attain or exceed the model quality of existing methods in both vision and language downstream tasks. We also conduct an in-depth analysis and discussion of the ScMoE architecture, investigating the proposed shortcut connection and exploring opportunities for further development.
@@
  • We propose the shortcut-connected MoE (ScMoE) architecture that breaks the conventional dependency between communication and computation in distributed MoE models, bypassing the restrictions imposed on current communication optimization techniques.
  • We develop an adaptive overlapping strategy for advancing expert parallelism with the shortcut-connected MoE, which significantly improves the efficiency of MoE models and ensures broad compatibility.
@@ 2.1 Sparsely-Gated Mixture of Experts
The sparsely-gated mixture-of-experts (Shazeer et al., 2017) (MoE) layer is composed of multiple multi-layer perceptron (MLP) sub-networks, termed “experts,” and employs a trainable gating network to selectively activate a subset of these experts during each iteration. Given $N$ expert networks $\{E_{i}\}_{1}^{N}$, gating network $G$ and input representation $x$, the output of MoE module can be written:
$MoE(x)=\sum_{i=1}^{N}G(x)_{i}E_{i}(x).$ (1)
@@
$G(x)=Softmax(\overline{TopK}(H(x),k)),$ (2)
@@
Leveraging sparse output of $G(x)$, this approach significantly increases the number of model parameters without causing a proportional increase in computational demand. The value of $k$ can be set to 1 or 2 or even higher values. Opting for a larger $k$ moves the model closer to the dense architecture, which generally results in higher prediction accuracy (Riquelme et al., 2021), but also leads to greater computational overhead.
@@
Figure 2 (a) illustrates the prevailing top-2 MoE architecture. Each Transformer block with MoE module, denoted by a light blue block and referred to as “Block-MoE,” replaces the MLP with a set of experts (“E1, E2, E3, E4”) and a gating network (“Gate”). Following prior work (Lepikhin et al., 2021; Du et al., 2022; Shen et al., 2023; Hwang et al., 2023; Lieber et al., 2025), the “Block-MoE” is interspersed with the conventional Transformer block, depicted as a gray block and referred to as “Block-MLP.” Additionally, various options exist for the frequency of MoE module placement, including placing an MoE module into every block (Jiang et al., 2024; Dai et al., 2024; Qwen, 2024; Databricks, 2024) or every four blocks (Zoph et al., 2022; Xue et al., 2024).
@@
$MoE(x)=SE(x)+\sum_{i=1}^{N}G(x)_{i}E_{i}(x).$ (6)
This method, originally proposed by DeepSpeed-MoE (Rajbhandari et al., 2022), activates the same number of experts as the standard MoE for computation while reducing dynamic expert selection and communication volume. Extensive empirical results from DeepSpeed-MoE and subsequent studies (Dai et al., 2024; Qwen, 2024) demonstrate that the shared expert architecture achieves model quality that is on par with or even superior to existing approaches, leading to its growing adoption (Xue et al., 2024; Wu et al., 2023; Chen et al., 2024; Gou et al., 2023; Gao et al., 2024).
@@ 2.2 Expert Parallelism
Figure 4: Illustrations of various ScMoE architectures with shortcut connections to different positions of the preceding layer: (a) “Pos-1” output, (b) “Pos-2” intermediate, and (c) “Pos-3” input. The red line indicates the transmission of the preceding-layer representations to the MoE via a shortcut connection. Details regarding pre-layer normalization and dropout procedures have been excluded for simplicity.
To facilitate efficient distributed training and inference of MoE models, expert parallelism is proposed to allocate unique experts to each distributed computing device such as GPU and TPU, and map tokens to their corresponding experts through All-to-All communication across participating devices (Lepikhin et al., 2021; He et al., 2021; Nie et al., 2022). As illustrated in Figure 3, the workflow of MoE employing expert parallelism is segmented into the following sequential operations: gate routing, input encode, All-to-All dispatch, expert computation, All-to-All combine, and output decode. To enhance the efficiency, input encode is employed to aggregate the token data layout to a contiguous format before All-to-All dispatch, and output decode is the inverse process after All-to-All combine. Furthermore, the integration of expert parallelism with other parallelisms (Hwang et al., 2023; Singh et al., 2023; Fedus et al., 2022; Zheng et al., 2022) has been explored to support the scaling of larger MoE models on extensive distributed systems. However, the All-to-All communication used for token transfer has been a primary bottleneck limiting the efficiency of distributed MoE models, as shown in Figure 1.
@@ 3 Shortcut-connected MoE Designs
In the prevailing Transformer-based model, the MoE module substitutes MLP to sequentially manipulate intermediate representations (Lepikhin et al., 2021; Du et al., 2022; Shen et al., 2023), impeding the efficacy of existing optimization strategies (He et al., 2022; Nie et al., 2022; Hwang et al., 2023; Zhang et al., 2023) due to the limited interaction within the MoE module. To address the aforementioned limitations, we propose the shortcut-connected MoE (ScMoE) architecture, which enables optimization opportunities for computation-communication overlap for expert parallelism.
@@ 3.1 Architectural Design
In this section, we introduce the shortcut-connected MoE (ScMoE) architecture. Unlike the prevailing MoE architectures, illustrated in Figure 2, which focus on processing intermediate representations within the current layer (the Transformer block containing the MoE), the ScMoE processes representations from both the current and preceding layers. Specifically, ScMoE employs a top-1 MoE module to handle representations from the preceding layer via a shortcut connection, while a shared expert processes the current-layer representations. These two operations are conducted independently, with their outcomes integrated into the final output of the current layer, facilitating communication and computation overlap between these two processes.
While the shared expert processes the same intermediate representations in the current layer as the prevailing MoE approaches, we explore three distinct preceding-layer representations for ScMoE’s top-1 MoE process, as illustrated in Figure 4. The configurations “Pos-1” (a), “Pos-2” (b), and “Pos-3” (c) represent shortcuts connecting different positions of the preceding layer: output, intermediate, and input, respectively. Given that $\mathcal{T}_{Atten}$, $\mathcal{T}_{SE}$, and $\mathcal{T}_{MLP}$ represent the durations of Attention, Shared Expert, and MLP, respectively, the corresponding overlap durations are (a) $\mathcal{T}_{Atten}+\mathcal{T}_{SE}$, (b) $\mathcal{T}_{Atten}+\mathcal{T}_{SE}+\mathcal{T}_{MLP}$, (c) $2\mathcal{T}_{Atten}+\mathcal{T}_{SE}+\mathcal{T}_{MLP}$.
Using the “Pos-2” configuration as an example, this ScMoE architecture can be formulated as follows:
@@
$+\sum_{i=1}^{N}G(\mathcal{H}^{MH}_{l})_{i}E_{i}(\mathcal{H}^{MH}_{l}),$
$\mathcal{H}^{MH}_{l+1}=\mathcal{H}^{MLP}_{l}+\text{MultiHead}^{(l+1)}_{MoE}(\mathcal{H}^{MLP}_{l}),$ (8)
@@
$\mathcal{H}^{MLP}_{l}=\mathcal{H}^{MH}_{l}+\text{MLP}^{(l)}(\mathcal{H}^{MH}_{l}),$ (9)
$\mathcal{H}^{MH}_{l}=\mathcal{H}_{l-1}+\text{MultiHead}^{(l)}_{MLP}(\mathcal{H}_{l-1}),$ (10)
where $\mathcal{H}_{l+1}^{\text{ScMoE}}$ refers to the output from the MoE sub-layer, $\mathcal{H}_{l+1}^{MH}$ signifies the output from the Multi-Head Attention (MultiHead) sub-layer $\text{MultiHead}^{(l+1)}_{MoE}(\cdot)$ in the $(l+1)$-th Transformer block (“Block-MoE”). $SE^{(l+1)}(\cdot)$ denotes the shared expert while $E_{1},...,E_{N}$ represent the $N$ gate-routed experts. The gating network $G(\cdot)$ is referred to as Equation 2. $\mathcal{H}_{l}^{MLP}$ and $\mathcal{H}_{l}^{MH}$ are the outputs of the MLP sub-layer $\text{MLP}^{(l)}(\cdot)$ and the MultiHead sub-layer $\text{MultiHead}^{(l)}_{MLP}(\cdot)$, respectively, in the $l$-th Transformer block (“Block-MLP”). Note that we omit the pre-layer normalization and dropout for simplicity.
In our experiments involving three shortcut-connected positions, models configured with “Pos-2” achieve the highest accuracy in both vision and language cases, while also ensuring substantial overlap duration. As a result, we favor selecting “Pos-2” for practical development. Moreover ,the “Pos-2” configuration is used to elucidate the overlapping strategy in Section 3.2. The specifics of the other two configurations can be inferred by analogy.
Additionally, our proposed ScMoE architecture can be adapted to support MoE models with varying MoE placement frequencies. As illustrated in Figure 5, the ScMoE architecture can be integrated into MoE models that incorporate an MoE module into every Transformer block. With more frequent MoE placement, the potential overlap duration for each MoE module is minimized, and the “Pos-1” configuration has already fully utilized the computation duration for the overlap. Conversely, less frequent MoE placement extends the potential overlap duration for each MoE module, which may lead to increased acceleration.
Figure 5: Illustration showcasing the application of the ScMoE (Pos-1) architecture to the MoE model, wherein the MoE module is integrated into each Transformer block.